# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v4

@@ A.6 Proof and Discussion of Theorem 4
Proof.
$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq. (LABEL:app_eq_11), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same. Thus we need to change the Eq. (LABEL:app_eq_12) as
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (29)
[…]
Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization. In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution. In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by Eq. (LABEL:app_ep_20) will be controlled.
A.7 Sparse Reward Task
@@ Appendix B Pseudo Code of TAD
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
@@
2:     //$Model\ Training$
3:     for update $\text{update step}=1,2,...,U$ do
@@
13:        for sample $\text{sample step}=1,2,...,S$ do
@@ C.1 Details of All Task Combinations
[…]
  • Cartpole-balance&sparse. This task combination includes two tasks: Cartpole-balance and Cartpole-balance__\__sparse, which both hope to balance an unactuated pole with dense and sparse rewards respectively. The optimal actions of these tasks are both hoped to balance the agent and thus TDR here is 0.
  • Walker-stand&walk&prostrate&flip. This task combination includes four tasks: Walker-stand, Walker-walk, Walker-prostrate, and Walker-flip. Walker-stand hopes the height of a two-leg robot to be larger than the target height. Walker-walk hopes the height of the improved planar walker to be larger than a target height and the speed of the robot to be larger than another target speed. Walker-prostrate hopes the height of the robot to be lower than the target height. Finally, Walker-flip hopes the robot to stand and move forward to the target speed by executing a rapid twist and jump. In detail, their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{stand}}=&\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\\ \mathcal{R}_{\text{prostrate}}=&\text{tolerance}(\text{height},(0.0,0.2)).\\ \mathcal{R}_{\text{flip}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{angmomentum},5,\infty).\end{split}$ (31)
@@
  • Cheetah-run&run__\__back&flip&flip__\__back. This task combination includes four tasks: Cheetah-run, Cheetah-run__\__back, Cheetah-flip, and Cheetah-flip__\__back. Cheetah-run hopes to control a running planar biped to run forward within a target speed. Cheetah-run__\__back, differently, hopes to control the Cheetah robot to run backward within a target speed. Cheetah-flip hopes the robot to move forward to the target speed by executing a rapid twist and jump. Similarly, Cheetah-flip__\__back hopes to control the robot to move backward by flipping.
@@
    Obviously, in this situation, for all states, the optimal actions of Cheetah-run and Cheetah-run__\__back are opposite, and TDR here is also huge.
Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
&balance__\__sparse &prostrate&flip &flip&flip__\__back
CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
CURL (w/ r) $\Pi_{3}$ 987.3 $\pm$ 12.9 265.0 $\pm$ 4.7 236.1 $\pm$ 5.8
SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
SAC+AE (w/ r) $\Pi_{3}$ 988.5 $\pm$ 7.0 257.2 $\pm$ 8.5 239.9 $\pm$ 12.4
Table 4: Performance (mean $\pm$ std) in DMC of CURL, CURL (w/ r), SAC+AE, and SAC+AE (w/ r).
Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run__\__mass Walker-walk__\__mass
@@
[…]
@@ C.2 Details of All Task Distributions
Now we introduce the three task distributions in our experiments, which are designed based on existing tasks in DeepMind control suite for testing the generalization of trained agents.
  • Cheetah__\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah__\__run in DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
  • Pendulum__\__angle($\alpha,\beta$). This task distribution is designed in this paper with parameter $-1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh__\__swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
    $\begin{split}\mathcal{R}_{\text{Pendulum}\_\text{angle}}(\alpha,\beta)=\text{tolerance}(\text{angle},(&\arccos\alpha,\\ &\arccos\beta)).\end{split}$ (34)
    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and test tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
  • Walker__\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker__\__run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. For each seed and each task setting, it will take around 3 days. As for the hyper-parameters, we follow previous works (Ha and Schmidhuber 2018; Hafner et al. 2019b, a) and select 2 as the action repeat for all experiments following (Hafner et al. 2019a).
@@ C.3 Visualization Results for Task Generalization
Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for test tasks), we plot the speed of the agent as a function of the timestep in Fig. 6. As depicted, for each task, the agent trained by TAD will quickly improve its speed until reaching the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information. Consequently, TAD is aware of different tasks and can successfully generalize to unseen test tasks. We also provide videos of these trajectories in supplementary materials.
@@
Figure 6: Visualization of the trained TAD agent in the task distribution of Cheetah__\__speed. We plot the speed of the agent as a function of timesteps in all tasks.

### v5

@@ A.6 Proof and discussion of Theorem 4
Proof A.7.
$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq. (20), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same. Thus, we need to change the Eq. (21) as
$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}-\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (30)
here $p(l)=\int p(\mathcal{M})d_{\tilde{\mathcal{M}},\pi}(l)d\mathcal{M}$, and $p(\mathcal{M}|l)=p(\mathcal{M})d_{\tilde{\mathcal{M}},\pi}(l)/p(l)$ is the posterior distribution.
Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization. In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution. In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by Eq. (29) will be controlled.
A.7 Sparse reward tasks
@@ Appendix B Pseudo code of TAD
0:$M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, NN test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
@@
[…]
3:  for $\text{update step}=1,2,...,U$ do
@@
13:   for $\text{sample step}=1,2,...,S$ do
@@ C.1 Details of all task combinations
In this part, we first roughly discuss the reward function of tasks in our experiments to better understand their TDR. These reward functions always defined by tolerance function in DeepMind control suite [8], which is a smooth function with parameters $\text{tolerance}(x,\text{bounds}=(\text{lower},\text{upper}))$ and hope the value of $x$ is within $(\text{lower},\text{upper})$. More details about tolerance function can be found in [8].
  • Cartpole-balance&sparse. This task combination includes two tasks: Cartpole-balance and Cartpole-balance_\_sparse, which both hope to balance an unactuated pole with dense and sparse rewards, respectively. The optimal actions of these tasks are both hoped to balance the agent, and thus, TDR here is 0.
  • Walker-stand&walk&prostrate&flip. This task combination includes four tasks: Walker-stand, Walker-walk, Walker-prostrate, and Walker-flip. Walker-stand hopes the height of a two-legged robot to be larger than the target height. Walker-walk hopes the height of the improved planar walker to be larger than a target height and the speed of the robot to be larger than another target speed. Walker-prostrate hopes the height of the robot to be lower than the target height. Finally, Walker-flip hopes the robot to stand and move forward to the target speed by executing a rapid twist and jump. In detail, their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{stand}}=&\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}=&\text{tolerance}(\text{height},(1.2,\infty))*\text{tolerance}(\text{speed},(1,\infty)).\\ \mathcal{R}_{\text{prostrate}}=&\text{tolerance}(\text{height},(0.0,0.2)).\\ \mathcal{R}_{\text{flip}}=&\text{tolerance}(\text{height},(1.2,\infty))*\text{tolerance}(\text{angmomentum},5,\infty).\end{split}$ (32)
@@
  • Cheetah-run&run_\_back&flip&flip_\_back. This task combination includes four tasks: Cheetah-run, Cheetah-run_\_back, Cheetah-flip, and Cheetah-flip_\_back. Cheetah-run hopes to control a running planar biped to run forward within a target speed. Cheetah-run_\_back, differently, hopes to control the Cheetah robot to run backward within a target speed. Cheetah-flip hopes the robot to move forward to the target speed by executing a rapid twist and jump. Similarly, Cheetah-flip_\_back hopes to control the robot to move backward by flipping.
@@
    Obviously, in this situation, for all states, the optimal actions of Cheetah-run and Cheetah-run_\_back are opposite, and TDR here is also huge.
Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run_\_mass Walker-walk_\_mass
@@
Table 4: Generalization performance (mean $\pm$ std) over different task distributions in image-based DMC of the best policy. Numbers greater than 95 $\%$ of the best performance for each environment are bold.
@@ C.2 Details of all task distributions
Now we introduce the three task distributions in our experiments, which are designed based on existing tasks in the DeepMind control suite for testing the generalization of trained agents.
  • Cheetah_\_speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah_\_run in the DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
  • Pendulum_\_angle($\alpha,\beta$). This task distribution is designed in this paper with parameter $1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh_\_swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
    $\begin{split}\mathcal{R}_{\text{Pendulum}\_\text{angle}}(\alpha,\beta)=\text{tolerance}(\text{angle},(&\arccos\alpha,\arccos\beta)).\end{split}$ (35)
    Training tasks are with parameters (0.95,0.9-0.95,-0.9), (0.85,0.8-0.85,-0.8), (0.8,0.75-0.8,-0.75), (0.7,0.65-0.7,-0.65) and test tasks are with parameters (0.9,0.85-0.9,-0.85), (0,75,0.7-0,75,-0.7).
  • Walker_\_speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker_\_run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. Each seed and each task setting will take around 3 days. As for the hyper-parameters, we follow previous works [3, 4, 5] and select 2 as the action repeat for all experiments following [5].
@@ C.3 Visualization results for task generalization
Moreover, for each task sampled from the task distribution Cheetah_\_speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for test tasks), we plot the speed of the agent as a function of the timestep in Fig. 4. As depicted, for each task, the agent trained by TAD will quickly improve its speed until it reaches the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information. Consequently, TAD is aware of different tasks and can successfully generalize to unseen test tasks. We also provide videos of these trajectories in the supplementary materials.
@@
Figure 4: Visualization of the trained TAD agent in the task distribution of Cheetah_\_speed. We plot the speed of the agent as a function of timesteps in all tasks.
C.4 Implementation details
TAD is implemented based on a widely adopted Dreamer repository https://github.com/yusukeurakami/dreamer-pytorch and we follow its default setups, including hyperparameters and network architecture. For the task model $p_{\theta}(\mathcal{M}|h_{t},s_{t})$, we choose to implement it with a simple 3-layer MLP, which maps current state $s_{t}$ and hidden state $s_{t}$ to a probability simplex $\triangle^{M-1}$ (assume we have $M$ tasks).
C.5 TDR analyses in DMControl experiments
We would like to illustrate the usage of TDR with a few examples in our experiments, including Walker-stand&\&walk and Walker-walk&\&prostrate.
To simplify the analyses, we set the reward of each state in Walker-stand $\mathcal{M}_{1}$ to 1 if the height of the robot is higher than 1.2 and 0 otherwise. In Walker-prostrate $\mathcal{M}_{2}$, we set the reward of each state to 1 if the height is lower than 0.2 and 0 otherwise. In Walker-walk, $\mathcal{M}_{3}$, we set the reward of each state to 1 if the height is higher than 1.2 and the speed is greater than 1.0; otherwise, the reward is 0.
We first estimate the optimal $Q^{*}$ function of these three tasks. For any state action pair $(s,a)$ in $\mathcal{M}_{1}$, if the height of the current state $s$ is higher than 1.2, the optimal policy is to maintain the height and we have $Q^{*}_{\mathcal{M}_{1}}=1+\gamma+\gamma^{2}+...=\frac{1}{1-\gamma}$. Otherwise, if the height of $s$ is lower than 1.2, the optimal policy is to make the agent stand in one step and maintain a height of at least 1.2. We can deduce that $Q^{*}_{\mathcal{M}_{1}}=\gamma+\gamma^{2}+...=\frac{\gamma}{1-\gamma}$. The analyses of $Q^{*}_{\mathcal{M}_{2}},Q^{*}_{\mathcal{M}_{3}}$ are similar and we conclude as follows:
[…]
Consider a task combination $\mathcal{T}_{1}$ that is distributed equally over Walker-stand and Walker-walk, i.e., $\mathbb{P}(\mathcal{M}_{1})=\mathbb{P}(\mathcal{M}_{3})=0.5$. If height of $s\geq 1.2$ and speed of $s\geq 1.0$, $Q^{*}_{\mathcal{M}_{1}}(s,a)=Q^{*}_{\mathcal{M}_{3}}(s,a)=\frac{1}{1-\gamma}$. Thus, we have:
$\begin{split}D_{\text{TDR}}(\mathcal{T}_{1},s)=&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[Q_{\mathcal{M}}^{*}(s,a)]=\frac{1}{1-\gamma}-\frac{1}{1-\gamma}=0.\end{split}$ (38)
Based on Theorem 3, for some policy $\pi\in\Pi_{1},\Pi_{2}$:
$J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}_{1}}\left[J_{\mathcal{M}}(\pi)\right]\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T}_{1},s)]=0.$ (39)
Consequently, in the task distribution $\mathcal{T}_{1}$, there is no gap between the optimal performance achievable by the policies in $\Pi_{1},\Pi_{2}$ and the actual optimal value. A TDR of 0 indicates that policies in $\Pi_{1}$ and $\Pi_{2}$ are sufficient to achieve optimal performance in $\mathcal{T}_{1}$..
In contrast, consider a task combination $\mathcal{T}_{2}$ that is distributed equally over Walker-walk and Walker-prostrate, i.e., $\mathbb{P}(\mathcal{M}_{2})=\mathbb{P}(\mathcal{M}_{3})=0.5$. For any $s$, $\frac{1}{2}(Q^{*}_{\mathcal{M}_{2}}(s,a)+Q^{*}_{\mathcal{M}_{3}}(s,a))\leq\frac{1}{2}(\frac{1}{1-\gamma}+\frac{\gamma}{1-\gamma})$. Thus, we have:
$\begin{split}D_{\text{TDR}}(\mathcal{T}_{2},s)=&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}_{2}}[Q_{\mathcal{M}}^{*}(s,a)]\geq\frac{1}{1-\gamma}-\frac{1}{2}\frac{1+\gamma}{1-\gamma}=0.5.\end{split}$ (40)
Consequently, based on Theorem 3, for any policy $\pi\in\Pi_{1},\Pi_{2}$:
$J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}_{2}}\left[J_{\mathcal{M}}(\pi)\right]\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T}_{2},s)]=\frac{1}{2}\frac{1}{1-\gamma}.$ (41)
In other words, our conclusion provides a non-trivial gap between the optimal performance achievable by the policies in $\Pi_{1},\Pi_{2}$ and the actual optimal value. A non-zero TDR shows that policies in $\Pi_{1},\Pi_{2}$ can only obtain half the optimal performance in $\mathcal{T}_{2}$.
Our experimental results in Table 5 also confirm our analysis regarding the effect of TDR. While policies in $\Pi_{1},\Pi_{2}$ perform appealingly in Walker-stand&\&walk, their performance degenerates significantly in Walker-walk&\&prostrate.
Algorithm Walker-stand&\&walk Walker-walk&\&prostrate
CURL 563.4 $\pm$ 16.0 441.0 $\pm$ 7.7
SAC+AE 683.8 $\pm$ 164.3 445.8 $\pm$ 2.6
PlaNet 990.6 $\pm$ 3.8 445.4 $\pm$ 5.1
Dreamer 992.3 $\pm$ 1.7 449.7 $\pm$ 12.5
TAD 994.6 $\pm$ 0.1 855.3 $\pm$ 56.2
Table 5: Performance (mean $\pm$ std) over different task combinations in DMC of the best policy. Numbers greater than 95 % of the best performance for each environment are bold.
C.6 MSE loss of imagination images in DMControl experiments
In Table 6, we have provided quantitative results on the models’ prediction error of imagined observations. As shown below, the prediction error of Dreamer gradually increases over the time horizon, while TAD exhibits no significant increasing trend. We attribute this phenomenon to TAD’s ability to access diverse trajectories sampled from different tasks during training.
Algorithm Task step 10 step 15 step 20 step 25 step 30 step 35
Dreamer Cheetah-run 19.2 18.8 28.7 29.3 30.1 32.5
Cheetah-flip 18.6 24.3 34.8 37.5 33.5 31.1
TAD (Ours) Cheetah-run 19.3 23.2 21.3 23.9 25.6 23.3
Cheetah-flip 17.9 18.8 21.1 19.5 17.3 19.4
Table 6: MSE Loss of imagination results of TAD and Dreamer
C.7 Additional results with task parameters
The core idea of TAD is to regularize the world model to distinguish tasks with $p_{\theta}(\mathcal{M}|h_{t},s_{t})$, with $\mathcal{M}$ being the task indicator. While the current implementation uses a set of discrete training tasks, we would like to clarify that our method is not limited to discrete spaces of tasks. For continuous spaces of tasks, when the task parameters are available, we can optimize TAD by directly predicting the continuous task parameters. When the task parameters are unavailable, we can switch to the discrete implementation, where we randomly sample $M$ tasks to construct a set of discrete training tasks and optimize TAD by predicting the task indices. Obviously, the task indices are less informative than task parameters, because they can only identify different tasks, while providing no information about the relationships or similarities among them. We have provided additional experiments of TAD with access to continuous task parameters in Table 7, labeled as TAD-Continuous. Results in the table indicate that our method using the discrete implementation (TAD-CE/SC) achieves performance comparable to its continuous counterpart.
Train&Test Train Test Train Test
RL2(zero-shot) $\Pi_{3}$ 1070.7 $\pm$ 109.7 -70.3 $\pm$ 6.7 191.9 $\pm$ 50.8
RL2(few-shot) $\Pi_{3}$ 1006.9 $\pm$ 26.4 -146.9 $\pm$ 0.4 268.8 $\pm$ 30.2
MAML(few-shot) 429.3 $\pm$ 81.4 -121.0 $\pm$ 37.1 205.3 $\pm$ 34.7
TAD-Continuous (Ours) $\Pi_{3}$ 1539.9 $\pm$ 102.0 -47.8 $\pm$ 1.1 -47.5 $\pm$ 1.3 302.9 $\pm$ 36.3 304.6 $\pm$ 39.0
[…]